# Blade: A package for block-triangular form improved Feynman integrals decomposition

### v1

@@ Acknowledgements
We thank Li-Hong Huang, Xiang Li, Bernhard Mistlberger, Tiziano Peraro, Ao Tan, Zi-Hao Wu, and Yang Zhang for many useful communications and discussions. We thank Huan-Yu Bi, Long Chen, Xiang Chen, Yuan-Hong Guo, Chuan-Qi He, Rui-Jun Huang, Xu-Hang Jiang, Xiang Li, Huai-Min Yu and Yu-Dong Zhang for providing numerous invaluable feedback on the package. The work was supported in part by the National Natural Science Foundation of China (Grants No. 12325503, No. 11975029), the National Key Research and Development Program of China under Contracts No. 2020YFA0406400, and the High-performance Computing Platform of Peking University. The research of XL was also supported by the ERC Starting Grant 804394 HipQCD and by the UK Science and Technology Facilities Council (STFC) under grant ST/T000864/1. JaxoDraw [86] was used to generate Feynman diagrams.
@@ B.1 Point group of Feynman integral
In Section II.3, we highlighted that linear transformations of loop momenta enable us to relate any given Feynman integral to its canonical form, which we will define next. The only condition we have imposed so far on the canonical form is that each loop momentum $l_{i}$ needs to serve as the propagating momentum of one of the propagators, as much as possible. We refer to the propagator with propagating momentum $l_{i}$ as the $i$-th loop base. However, there can be different choices of loop bases, and even for a given choice of loop bases, there can be different ways to label each propagating momentum. Starting from a candidate for the canonical form, which means a choice of loop bases has already been made, we can generate many more candidates by applying the following operations:

### v2

@@ Acknowledgements
We thank Li-Hong Huang, Xiang Li, Bernhard Mistlberger, Tiziano Peraro, Ao Tan, Zi-Hao Wu, and Yang Zhang for many useful communications and discussions. We thank Huan-Yu Bi, Long Chen, Xiang Chen, Yuan-Hong Guo, Chuan-Qi He, Rui-Jun Huang, Xu-Hang Jiang, Xiang Li, Huai-Min Yu and Yu-Dong Zhang for providing numerous invaluable feedback on the package. The work was supported in part by the National Natural Science Foundation of China (Grants No. 12325503, No. 11975029), the National Key Research and Development Program of China under Contracts No. 2020YFA0406400, and the High-performance Computing Platform of Peking University. The research of XG was also supported by the United States Department of Energy, Contract DE-AC02-76SF00515. The research of XL was also supported by the ERC Starting Grant 804394 HipQCD and by the UK Science and Technology Facilities Council (STFC) under grant ST/T000864/1. JaxoDraw [89] was used to generate Feynman diagrams.
@@ B.1 Point group of Feynman integral
In Section 2.3, we highlighted that linear transformations of loop momenta enable us to relate any given Feynman integral to its canonical form, which we will define next. The only condition we have imposed so far on the canonical form is that each loop momentum $l_{i}$ needs to serve as the propagating momentum of one of the propagators, as much as possible. We refer to the propagator with propagating momentum $l_{i}$ as the $i$-th loop base. However, there can be different choices of loop bases, and even for a given choice of loop bases, there can be different ways to label each propagating momentum. Starting from a candidate for the canonical form, which means a choice of loop bases has already been made, we can generate many more candidates by applying the following operations:
  • 1.